AnimeZodiac:

I'm trying to figure out a math problem and it's on multiping exponents can someone help?

toga:

i can type it in

AnimeZodiac:

this is an example : (4x-5)(6-1x)

toga:

To multiply the two expressions (4x - 5) and (6 - x), we need to use the FOIL method which stands for First, Outer, Inner, and Last. First, we multiply the first terms in each expression: 4x * 6 = 24x Next, we multiply the outer terms: 4x * (-x) = -4x^2 Then, we multiply the inner terms: (-5) * 6 = -30 Finally, we multiply the last terms in each expression: (-5) * (-x) = 5x Now, we can combine all the terms we just found: 24x - 4x^2 - 30 + 5x Simplifying the above expression further: -4x^2 + 29x - 30 Therefore, the product of (4x - 5) and (6 - x) is -4x^2 + 29x - 30.

AnimeZodiac:

please explain more I'm in special ed if you don't know what it is just tell me

AnimeZodiac:

sry

toga:

To multiply two expressions, we need to take each term of one expression and multiply it by each term of the other expression. In this case, we have two expressions, (4x-5) and (6-x). To multiply these expressions using the FOIL method, we first need to multiply the First terms of each expression. The first terms here are 4x and 6, so we multiply them to get 24x. Next, we multiply the Outer terms, which are 4x and -x. We get -4x^2 as a result. Then, we multiply the Inner terms, which are -5 and 6. This gives us -30. Finally, we multiply the Last terms, which are -5 and -x. We get 5x. Now, we can combine all the terms we just found: 24x - 4x^2 - 30 + 5x Simplifying the above expression further: -4x^2 + 29x - 30 So the product of (4x - 5) and (6 - x) is -4x^2 + 29x - 30.

toga:

@animezodiac wrote:
sry
it is ok

toga:

@animezodiac wrote:
please explain more I'm in special ed if you don't know what it is just tell me
Everyone has something unique about them. For me, it's OCD, and that's okay do not be sorry

AnimeZodiac:

@toga wrote:
@animezodiac wrote:
please explain more I'm in special ed if you don't know what it is just tell me
Everyone has something unique about them. For me, it's OCD, and that's okay do not be sorry
thx and what is OCD

toga:

@animezodiac wrote:
@toga wrote:
@animezodiac wrote:
please explain more I'm in special ed if you don't know what it is just tell me
Everyone has something unique about them. For me, it's OCD, and that's okay do not be sorry
thx and what is OCD
OCD stands for Obsessive-Compulsive Disorder. It is a mental health condition that is characterized by repeated unwanted thoughts, images, or urges (obsessions) and repetitive behaviors or mental acts (compulsions) that are performed in response to these obsessions. People with OCD may feel compelled to perform these rituals in order to alleviate anxiety or prevent harm to themselves or others, even though they may have no logical connection to the situation. OCD can significantly interfere with daily life and relationships, but it is treatable with therapy and medication.

AnimeZodiac:

ok

AnimeZodiac:

thx

toga:

@animezodiac wrote:
thx
np

AnimeZodiac:

and sry but im a girl and is 17 years old you

toga:

@animezodiac wrote:
and sry but im a girl and is 17 years old you
I am a girl and I am almost 15

AnimeZodiac:

oh my friend Carbohol that's his last name and he just turned 16 he's a boy a cute one at that

AnimeZodiac:

dont now how to spell his name

toga:

can we talk in DMS

AnimeZodiac:

?

AnimeZodiac:

what is white mean

AnimeZodiac:

and what is DMS

toga:

Private message

AnimeZodiac:

ohh